if the electric field is 100N/C at a distance of 50 cm from a point charge is the value of q? 


1
Expert's answer
2021-05-30T13:29:10-0400

By definition, the electric field of a point charge is given as follows:


"E = k\dfrac{q}{r^2}"

where "q" is the charge, "r = 50cm = 0.5m" is the distance from the charge, "k = 9\times 10^9N\cdot m^2/C^2" is the Coulomb's constant. Substituting "E = 100N\/C" and expressing "q", obtain:


"q = \dfrac{Er^2}{k} = \dfrac{100\cdot 0.5^2}{9\times 10^9} \approx 2.8\times 10^{-9}C"

Answer. "2.8\times 10^{-9}C".